# Trap Identification in Molecular Charge Transport Networks

### v1

@@ 4 Results
In the special case with only a single trap $N_{t}=1$ the set $Q(1)$ is given by a single node index $i$, and we can inspect $\tau(1)/\tau^{i}(1)$. This is shown for all $i$ for the simulated BCP system in Figure 2. It is clear that there is a single molecule node whose presence in the molecular charge transport network affects the ToF by four orders of magnitude. While it is intuitive to call this particular node a trap in this specific case, it is less obvious in general and the precise definition of a threshold value here is hardly possible. Another drawback of using the inspection of the ratio $\tau(1)/\tau^{i}(1)$, or its more general form, is the need for many cumbersome ToF calculations of the (${N\choose N_{t}}$ different scenarios. This shows that a general method that identifies traps in molecular charge transport networks based on the graph structure alone without the need to actually calculate the dynamical properties is of great usefulness.
4.1 Identification of Single-molecule Trap
@@
4.1.1 Results from Graph Decomposition Method
@@
Before turning to the spectral clustering method as proposed in Section 3, we first show the results of the graph-theoretical decomposition method [22], whose technical details are summarized in B. The method contains three adjustable parameters, $\alpha,\beta,\gamma$. The first two are related to details of the connectivity of the graph, whereas $\gamma$ is used to define a threshold involving the ratios of transition rates. In application to multi-node traps or trap regions, a choice of $\alpha,\beta=0.02$ and $\gamma=0.2$ has been reported before [21].
@@
4.1.2 Results from Spectral Clustering Method
@@
4.1.3 Relation between Cost function and ToF
@@
4.2 Identification of Multiple Distributed Traps
@@
The previous section shows that the spectral clustering method can reliably identify a single trap node as it occurs in the modeled BCP system. In the following, we will scrutinize if the same holds in a system with multiple trap molecules which are not connected with each other – a scenario we refer to as multiple distributed traps. Within the spectral clustering method, we expect that having multiple distributed traps will increase the cluster number $k$ for the K-means clustering steps as given by Algorithm 1. As the simulated BCP system does not have multiple distributed traps, we modify it taking the trap node we found and 4 other nodes ($Q=\{v_{1}=v_{\mathrm{trap}},v_{2}\cdots,v_{5}\}$), which are not connected by an edge in the charge transport network and set their site energies all to $E=-1.8\,\mathrm{eV}$, which is close to the site energy of the original trap node. It can be seen from the dependence of the ratio $\tau(1)/\tau^{Q}(1)$ on $E$ in Figure 6(a) that such a value points indeed to a very pronounced tapping effect of the charge carrier.
@@
Turning to our spectral clustering method, we see in Figure 7(a) and (b) that the largest value for the ratio is still at $i=2$, as was the case for the single trap situation. However, unlike that situation, when inspecting the first two normalized eigenvectors in Figure 7(d,e), one can clearly see five and four relevant non-zero elements, respectively. Overall, the determined number of unique induced subgraphs is $k=6$, which indeed implies that there are 5 traps in the system shown in Figure 7(f): the five individual distributed trap nodes and one cluster that is the rest of the system. This shows the necessity of the addition step in Algorithm 1. The result $k=6$ is also relatively robustly with respect to the choice of the method’s parameter $a$ as shown in Figure 7(c). Only when $a>0.99$ is used, the spectral clustering method yields $k=3$. So again, using Algorithm 2 with $a=0.9$ finds the correct number of clusters $k=6$, and also the right trap nodes.
4.3 Identification of Trap Region
@@
Figure 8(b) shows the result of the GD method to identify the traps: the total number of clusters $n_{\text{cl}}$ and the maximum cluster size $n_{m}$ when $m$ is one of the nodes in $Q_{\text{region}}$ for different values of $\gamma$. When $\gamma=10^{-6}$, $n_{m}=5$, and $Q_{\text{region}}$ is successfully identified as a single cluster and the rest of the 995 nodes as another cluster. As in Section 4.1, the isolated trap node is not identified as a separate cluster. As $\gamma$ increases, more clusters appear, and when $\gamma>10^{-2}$, the five low-energy nodes in $Q_{\text{region}}$ are no longer partitioned into one cluster.
@@
5 Conclusion
Traps play an important role in defining the electrical properties of organic semiconductors and their functionality electronic devices. They significantly influence the charge transport mechanism in OSCs, and allow for a tuning of the charge mobility as indicated by variations in the time-of-flight (ToF) measurements. Understanding and eventual control of the charge mobility is often sought via multiscale modeling of the materials which leads to its representation as a molecular charge transport network. In this study, we have developed and implemented a novel method based on spectral clustering to identify traps within such molecular charge transport network of multiscale-modeled OSCs. Our approach avoids the complexities associated with calculating multiple-carrier ToF and overcomes the need for system-dependent parameters, making it a robust method for general multiscale model system. Leveraging the eigenvalue ratio and eigenvector entries allows the application of K-means clustering on the eigenvector elements from the Laplacian matrix, reinforcing the robustness of our spectral clustering approach.
The effectiveness of our method is demonstrated through the analysis of the multiscale-model BCP device, where we successfully identified a single-molecule trap, distributed traps, and region of traps coexisting with single-molecule traps. This capability to identify and quantify traps without extensive parameter tuning marks a significant advancement over other candidate methods such as the watershed algorithm and graph-theoretic decomposition. Our proposed method and the findings from this study can potentially enhance our understanding of trap dynamics in OSCs and also offer a powerful tool for optimizing the design and performance of these materials in practical applications.
@@ Appendix B Graph Decomposition Method
The full theory and implementation details of the Graph Decomposition (GD) method are in [22]. Here a summary of the GD method and its parameters is provided. The GD algorithm begins by taking a vertex of minimal vertex degree in $\mathbf{G}$ and uses this vertex as the basis of a cluster $\bar{\mathbf{G}}$. Then each node $v^{\prime}$ adjacent to $v$ is checked, and added to $\bar{\mathbf{G}}$ if certain criteria are satisfied. The process is repeated until no more nodes can be added to $\bar{\mathbf{G}}$. At that stage, the nodes of subgraph $\bar{\mathbf{G}}$ are removed from $\mathbf{G}$ and classified as a cluster. The algorithm begins again until all nodes are classified. The criteria for including a node into a cluster are:

### v2

@@ IV Results
In the special case with only a single trap $N_{t}=1$ the set $Q(1)$ is given by a single node index $i$, and we can inspect $\tau(1)/\tau^{i}(1)$. This is shown for all $i$ for the simulated BCP system in Figure 2. It is clear that there is a single molecule node whose presence in the molecular charge transport network affects the ToF by four orders of magnitude. While it is intuitive to call this particular node a trap in this specific case, it is less obvious in general and the precise definition of a threshold value here is hardly possible. Another drawback of using the inspection of the ratio $\tau(1)/\tau^{i}(1)$, or its more general form, is the need for many cumbersome ToF calculations of the (${N\choose N_{t}}$ different scenarios. It also fails to reveal any details about the physical nature of the trap(s). This shows that a general method that identifies traps in molecular charge transport networks based on the graph structure alone without the need to actually calculate the dynamical properties is of great usefulness. Its microscopic insight must, however, be compatible with the effects of traps on macroscopic observables such as the ToF here. Finally, we emphasize that due to the modifications of the BCP baseline system, and some modeling choices in the multiscale model, the reported charge transport properties, such as time-of-flight or mobilities, should not be compared to actual measured material properties, if available.
IV.1 Identification of Single-molecule Trap
@@
IV.1.1 Results from Graph Decomposition Method
@@
Before turning to the spectral clustering method as proposed in Section III, we first show the results of the graph-theoretical decomposition method [24], whose technical details are summarized in B. The method contains three adjustable parameters, $\alpha,\beta,\gamma$. The first two are related to details of the connectivity of the graph, whereas $\gamma$ is used to define a threshold involving the ratios of transition rates. In application to multi-node traps or trap regions, a choice of $\alpha,\beta=0.02$ and $\gamma=0.2$ has been reported before [23].
@@
IV.1.2 Results from Spectral Clustering Method
@@
IV.1.3 Relation between Cost function and ToF
@@
IV.1.4 Effect of an Applied Electric Field
Figure 6: (a) The 1-carrier ToF $\tau$, and the ToF $\tau^{v_{\text{trap}}}$ when removing the node $v_{\text{trap}}$ as a function of the electric field.
To further validate the robustness of our spectral clustering method, we investigate its performance under non-equilibrium conditions by introducing an external electric field as typically used in operating conditions for devices based on organic semiconductors. The charge dynamic is then a drift-diffusion process rather than a diffusion. The electric field we applied is $\vec{F}=[F_{x},0,0]$ where $F_{x}$ has values in the range 00 to $10^{8}\,\mathrm{V/m}$.
The inclusion of an electric field in the Marcus rates modifies the transition rates between molecules, as the field introduces a directional bias in charge carrier motion. Figure 6 shows that as the electric field strength increases above $10^{6}\,\mathrm{V/m}$, the ToF decreases significantly, indicating a faster charge transport along the direction of $\vec{F}$. Using the proposed method, the node $v_{\text{trap}}$ is found as a single cluster across the range of the studied electric field. Removing $v_{\text{trap}}$ results in a significant decrease in ToF as indicated by the gap between $\tau$ and $\tau^{v_{\text{trap}}}$. As the electric field $|\vec{F}|>10^{6}\,\mathrm{V/m}$, the gap between $\tau$ and $\tau^{v_{\text{trap}}}$ shows signs of reducing. This phenomenon suggests that an electric field, having the direction drift force for the charge dynamics, can potentially reduce the trap effects.
IV.2 Identification of Multiple Distributed Traps
@@
The previous section shows that the spectral clustering method can reliably identify a single trap node as it occurs in the modeled BCP system. In the following, we will scrutinize if the same holds in a system with multiple trap molecules which are not connected with each other – a scenario we refer to as multiple distributed traps. Within the spectral clustering method, we expect that having multiple distributed traps will increase the cluster number $k$ for the K-means clustering steps as given by Algorithm 1. As the simulated BCP system does not have multiple distributed traps, we modify it taking the trap node we found and 4 other nodes ($Q=\{v_{1}=v_{\mathrm{trap}},v_{2}\cdots,v_{5}\}$), which are not connected by an edge in the charge transport network and set their site energies all to $E=-1.8\,\mathrm{eV}$, which is close to the site energy of the original trap node. It can be seen from the dependence of the ratio $\tau(1)/\tau^{Q}(1)$ on $E$ in Figure 7(a) that such a value points indeed to a very pronounced trapping effect of the charge carrier.
@@
Turning to our spectral clustering method, we see in Figure 8(a) and (b) that the largest value for the ratio is still at $i=2$, as was the case for the single trap situation. However, unlike that situation, when inspecting the first two normalized eigenvectors in Figure 8(d,e), one can clearly see five and four relevant non-zero elements, respectively. Overall, the determined number of unique induced subgraphs is $k=6$, which indeed implies that there are 5 traps in the system shown in Figure 8(f): the five individual distributed trap nodes and one cluster that is the rest of the system. This shows the necessity of the addition step in Algorithm 1. The result $k=6$ is also relatively robustly with respect to the choice of the method’s parameter $a$ as shown in Figure 8(c). Only when $a>0.9$ is used, the spectral clustering method yields $k=2$. So again, using Algorithm 2 with $a=0.9$ finds the correct number of clusters $k=6$, and also the right trap nodes.
IV.3 Identification of Trap Region
@@
Figure 9(b) shows the result of the GD method to identify the traps: the total number of clusters $n_{\text{cl}}$ and the maximum cluster size $n_{m}$ when $m$ is one of the nodes in $Q_{\text{region}}$ for different values of $\gamma$. When $\gamma=10^{-6}$, $n_{m}=5$, and $Q_{\text{region}}$ is successfully identified as a single cluster and the rest of the 995 nodes as another cluster. As in Section IV.1, the isolated trap node is not identified as a separate cluster. As $\gamma$ increases, more clusters appear, and when $\gamma>10^{-2}$, the five low-energy nodes in $Q_{\text{region}}$ are no longer partitioned into one cluster.
@@
Figure 11: Visualization of two non-energy based trapping scenarios: (a) the node $v_{\text{trap}}$ (blue) with the lowest energy and its (green) represent different molecular types with non-symmetric reorganization energies is correctly identfied as not being a trap. Two BCP molecules in red color, $v_{1}$ and $v_{2}$, are two single clusters identified as actual traps using the K-means clustering with $k=3$. (b) The neighbors of $v_{\text{trap}}$ are colored in light-dark and their electronic coupling element with $v_{\text{trap}}$ are small. Two BCP molecules in red color, $v_{1}$ and $v_{3}$, are two single traps identified using the K-means clustering with $k=3$ in this topology-based scenario.
IV.4 Non-symmetric reorganization energies
In materials containing different types of molecules and complex molecular geometries, the reorganization energy $R_{ij}$ for charge transfer between molecules is non-symmetric, i.e., $R_{ij}\neq R_{ji}$. In such cases, the rates for this transfer do not obey detailed balance and the equilibrium Boltzmann population argument mentioned in the Introduction cannot be applied, and the lowest energy site may not necessarily act as a trap, even if it has a significantly lower site energy compared to the rest of the network.
To illustrate that our proposed method is also able to correctly identify traps in such a situation, we modify the BCP system by introducing a non-symmetric reorganization energy for specific pairs of molecules. We select the node $v_{\text{trap}}$ with the lowest site energy $E_{v_{\text{trap}}}=-1.80\,\mathrm{eV}$ and adjust the reorganization energy $R$ such that $R_{v_{\text{trap}}\,j}=0.1\,\mathrm{eV}$ and $R_{j\,v_{\text{trap}}}=0.2\,\mathrm{eV}$. We show this setup in Figure 11(a), where the node $v_{\text{trap}}$ is colored in blue and its neighbors are in lighter color.
Due to the non-symmetric reorganization energy, the transition rate from $v_{\text{trap}}$ to its neighbors is not small, despite the low energy. This results in a situation where the low-energy site is not a trap because charge carriers can easily escape from it due to the high forward transition rate. In contrast, the spectral clustering method (see detailed results in Fig. 14 in the Appendix), which considers the full graph structure and transition rates, correctly identifies that this site is not a trap. Instead, the method identifies nodes $v_{1}$ and $v_{2}$ with site energy $E_{v_{1}}=E_{v_{2}}=-1.70\,\mathrm{eV}$ (in red color shown in Fig. 11(a)) as the isolated single traps in this system, putting those two nodes in their own cluster and the other 998 nodes in another.
The actual calculated ToF support the results of the spectral clustering method. For the system without removing any node we find $\tau=5.7\cdot 10^{-1}\,\mathrm{s}$. Removing $v_{\text{trap}}$, one obtains $\tau^{v_{\text{trap}}}=5.5\cdot 10^{-1}\,\mathrm{s}$, while removing $v_{1},v_{2}$ yields $\tau^{(v_{1},v_{2})}=1.4\times 10^{-3}\,\mathrm{s}$. We clearly see that removing the lowest energy site $v_{\text{trap}}$ has hardly any effect on the ToF, while the nodes identified from our proposed method influence the ToF by more than two orders of magnitude.
IV.5 Topological coupling-based traps
Next to mixed molecular material compositions, in general structural features can also lead to charge trapping behavior in materials. To mimic such a scenario within the as-modeled BCP baseline system, we modify the electronic coupling element $J_{v_{\text{trap}}\,j}$ of the lowest-energy site $v_{\text{trap}}$ with its neighboring sites $j$. Specifically, the electronic coupling elements $J_{v_{\text{trap}}\,j}$ between the lowest energy site $v_{\text{trap}}$ (highlighted in blue in Fig. 15) and its neighbors $j$ has range: $10^{-9}\,\mathrm{eV^{2}}>J_{v_{\text{trap}}\,j}^{2}>10^{-14}\,\mathrm{eV^{2}}$. Such a system is visualized in Figure 11(b), where the lowest energy site $v_{\text{trap}}$ is in color blue and its neighbors are in light-dark color.
In such a case, even if a site has a very low energy, the probability of charge carriers transitioning to or from this site is extremely low, resulting in low occupancy and hence no trapping effect in the non-steady state. While using a solely energy-based argument would still classify $v_{\text{trap}}$ as a trap due to its low energy, it does not exhibit any trapping effect in practice, as carriers are unlikely to be captured or released from this site. The ToF of this system is $\tau=6.0\cdot 10^{-1}\,\mathrm{s}$, which is not chnaged after removing $v_{\text{trap}}$. Applying our proposed spectral clustering method (see the full results in Fig. 15 in the Appendix), gives the same $v_{1}$ as before and a new site $v_{3}$ as two single clusters while the remaining 998 nodes form a single large cluster. Removing $v_{1},v_{3}$, the ToF is $\tau^{(v_{1},v_{3})}=6.3\cdot 10^{-2}\,\mathrm{s}$, indicating the real trapping effect of $v_{1},v_{3}$.
V Conclusion
In this study, we have developed and implemented a novel method based on spectral clustering to identify traps of widely different characteristics within molecular charge transport networks. Our approach is based on the application of K-means clustering on the eigenvector elements from the Laplacian matrix with a proposed heuristic based on eigenvalue ratio and eigenvector entries. The method avoids the complexities associated with calculating multiple-carrier ToF and overcomes the need for system-dependent parameters. Its effectiveness is demonstrated through the analysis of several trap types prepared with a baseline multiscale model. We successfully identified energy-based traps with a single-molecule, distributed molecule, and region characteristics. Our method also yields the correct identification of topology-based traps and is applicable to other, more general, situations (e.g., in mixtures with non-symmetric reorganization energies) in which the transition rates do not obey detailed balance.
This capability to identify and quantify traps without extensive parameter tuning marks a significant advancement over other candidate methods such as the watershed algorithm and graph-theoretic decomposition, or purely energy-based arguments. The added value of the spectral clustering method lies in its ability to capture the interplay between site energies, transition rates, and the overall network structure. This makes it a more robust and general approach for identifying traps in organic semiconductors, particularly in systems with complex charge transport dynamics.
@@ Appendix B Graph Decomposition Method
The full theory and implementation details of the Graph Decomposition (GD) method are in [24]. Here a summary of the GD method and its parameters is provided. The GD algorithm begins by taking a vertex of minimal vertex degree in $\mathbf{G}$ and uses this vertex as the basis of a cluster. Then each node $v^{\prime}$ adjacent to $v$ is checked, and added to $\bar{\mathbf{G}}$ if certain criteria are satisfied. The process is repeated until no more nodes can be added to $\bar{\mathbf{G}}$. At that stage, the nodes of subgraph $\bar{\mathbf{G}}$ are removed from $\mathbf{G}$ and classified as a cluster. The algorithm begins again until all nodes are classified.
The criteria for including a node into a cluster are:
@@
Appendix C Spectral Clustering for non-energy based traps
Figure 14: Trap identification results using spectral clustering methods on the BCP system where the site $v_{\text{trap}}$ and its neighbors have nonsymmetric reorganization energy. (a) The first ten eigenvalues of $L_{\text{rw}}$. (b) $\lambda_{i+1}/{\lambda_{i}}$ as a function of $i$. (c)-(d) The first and second normalized eigenvector elements of $L_{\text{rw}}$ as a function of node indices.
Figure 15: Trap identification results using spectral clustering methods on the BCP system where the node $v_{\text{trap}}$ and its neighbors have weak coupling elements. (a) The first ten eigenvalues of $L_{\text{rw}}$. (b) $\lambda_{i+1}/{\lambda_{i}}$ as a function of $i$. (c)-(d) The first and second normalized eigenvector elements of $L_{\text{rw}}$ as a function of node indices.